# Edge Zeta Functions and Eigenvalues for Spherical Buildings: A Uniform Hecke Algebra Approach

### v2

@@ 4.3.3 Closed formulas
$Z_{C}(X_{2}(\mathcal{B}),u)=\prod_{\chi:\,n_{\chi}\neq 0}\ \prod_{\zeta^{\,d}=1}\bigl(1-\zeta\,q^{(f_{\chi}-2\ell(w_{I}))/d}\,u^{\,c}\bigr)^{-m_{C,\chi}(\zeta)\,d_{\chi}}.$
@@
$Z_{C}(X_{2}(\mathcal{B}),u)=\prod_{\chi:\,n_{\chi}\neq 0}\bigl(1-(-1)^{|\mu|}\,q^{(f_{\chi}-2\ell(w_{I}))/2}\,u^{4}\bigr)^{-d_{\chi}n_{\chi}}.$
@@ 4.4 Tables of cycles and half-cycle length $m$
Some computations were carried out in CHEVIE for GAP3; for exceptional types we follow the simple–reflection labels used in [17, p. 1592, §85]. In types $\mathbf{B}/\mathbf{C}/\mathbf{D}$ we place the “branching” at the far end.
@@
(1) Higher zeta functions.
The edge zeta function considered here encodes geodesics in the 11–skeleton of the building. A next step is to define and study higher–dimensional zeta functions that count closed geodesics in higher skeleton of the building. The top–dimensional case would be the gallery zeta function, enumerating closed chamber galleries and generalizing the edge zeta function considered here. Such higher zeta functions should encode finer representation–theoretic and geometric information, potentially connecting to the cohomology of the building and to higher–rank analogues of Ihara’s determinant formulas.
(2) Regular elements and character values.
In Section 4.3.2, the quantities $\chi^{\lambda,\mu}(ue_{W_{I}})$ in type $\mathbf{C}$ (and their analogues $\chi^{\lambda}(ue_{S_{\mu}})$ in [18, §8.3]) take values in $\{0,\pm 1\}$. This pattern resembles the character values of Coxeter elements or some regular elements in Weyl groups [20].

### v3

@@ 4.3.3 Closed formulas
$Z_{c}(X_{2}(\mathcal{B})|_{C},u)=\prod_{\chi:\,n_{\chi}\neq 0}\ \prod_{\zeta^{\,d}=1}\bigl(1-\zeta\,q^{(f_{\chi}-2\ell(w_{I}))/d}\,u^{\,c}\bigr)^{-m_{C,\chi}(\zeta)\,d_{\chi}}.$
@@
$Z_{c}(X_{2}(\mathcal{B})|_{C},u)=\prod_{\chi:\,n_{\chi}\neq 0}\bigl(1-(-1)^{|\mu|}\,q^{(f_{\chi}-2\ell(w_{I}))/2}\,u^{4}\bigr)^{-d_{\chi}n_{\chi}}.$
@@ 4.4 Tables of cycles and half-cycle length $m$
Some computations were carried out in CHEVIE for GAP3; for exceptional types we follow the simple–reflection labels used in [undefp, p. 1592, §85]. In types $\mathbf{B}/\mathbf{C}/\mathbf{D}$ we place the “branching” at the far end.
@@
(1) Higher zeta functions
The edge zeta function considered here encodes geodesics in the 11–skeleton of the building. A next step is to define and study higher–dimensional zeta functions that count closed geodesics in higher skeleton of the building. At the top dimension, one obtains the gallery zeta function, enumerating closed chamber galleries and generalizing the edge zeta function considered here. Such higher zeta functions should encode finer representation–theoretic and geometric information, potentially connecting to the cohomology of the building and to higher–rank analogues of Ihara’s determinant formulas.
(2) Regular elements and character values
In Section 4.3.2, the quantities $\chi^{\lambda,\mu}(ue_{W_{I}})$ in type $\mathbf{C}$ (and their analogues $\chi^{\lambda}(ue_{S_{\mu}})$ in [undefr, §8.3]) take values in $\{0,\pm 1\}$. This pattern resembles the character values of Coxeter elements or some regular elements in Weyl groups [undeft].
@@
Acknowledgements.
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358.